1.) A stock just paid a dividend of $1.37. The dividend is expected to grow at 29.31% for three years and then grow at 3.42% thereafter. The required return on the stock is 11.32%. What is the value of the stock?

2.) A stock just paid a dividend of $1.98. The dividend is expected to grow at 25.37% for five years and then grow at 4.00% thereafter. The required return on the stock is 10.43%. What is the value of the stock?

Thank you so much!
Answer #1

1)

Year 1 dividend = 1.37 (1 + 0.2931) = 1.77155

Year 2 dividend = 1.77155 (1 + 0.2931) = 2.29079

Year 3 dividend = 2.29079 (1 + 0.2931) = 2.96222

Year 4 dividend = 2.96222 (1 + 0.0342) = 3.06353

Value at year 3 = D4 / required rate - growth rate

Value at year 3 = 3.06353 / 0.1132 - 0.0342

Value at year 3 = 3.06353 / 0.079

Value at year 3 = $38.77886

Value of stock = 1.77155 / (1 + 0.1132)1 + 2.29079 / (1 + 0.1132)2 + 2.96222 / (1 + 0.1132)3 + 38.77886 / (1 + 0.1132)3

Value of stock = $33.70

2)

Year 1 dividend = 1.98 (1 + 0.2537) = 2.48233

Year 2 dividend = 2.48233 (1 + 0.2537) = 3.11209

Year 3 dividend = 3.11209 (1 + 0.2537) = 3.90163

Year 4 dividend = 3.90163 (1 + 0.2537) = 4.89147

Year 5 dividend = 4.89147 (1 + 0.2537) = 6.13244

Year 6 dividend = 6.13244 (1 + 0.04) = 6.37774

Value at year 5 = D6 / required rate - growth rate

Value at year 5 = 6.37774 / 0.1043 - 0.04

Value at year 5 = 6.37774 / 0.0643

Value at year 5 = 99.18725

Value of stock = 2.48233 / (1 + 0.1043)1 + 3.11209 / (1 + 0.1043)2 + 3.90163 / (1 + 0.1043)3 + 4.89147 / (1 + 0.1043)4 + 6.13244 / (1 + 0.1043)5 + 99.18725 / (1 + 0.1043)5

Value of stock = $75.12